OpenStudy (s117):

Fred is playing tennis. For every 2 of his serves that land in, 4 serves land out. If he hit 28 serves in, how many serves landed out? 224 56 8 112

jagr2713 (jagr2713):

ok for every 2 servers 4 land out. so if he gets 4 serves 8 land out so he gt 28 serves. So 28*4=??

OpenStudy (s117):

112

jagr2713 (jagr2713):

correct. but i am a little confused so let me ask someone @pooja195 @freckles

pooja195 (pooja195):

seems rigth

OpenStudy (freckles):

\[\frac{\text{ serves in } }{ \text{ serves out } } =\frac{\text{ a given serves in number }} {\text{ what would be the out number for that given in number }} \\ \frac{2}{4}=\frac{28}{x } \\ \text{ reduce right fraction } \\ \frac{1}{2}=\frac{28}{x} \\ \text{ flip both sides } \\ \frac{2}{1}=\frac{x}{28} \\ 2=\frac{x}{28} \text{ since } \frac{2}{1}=2 \\ x=2(28)\]

OpenStudy (freckles):

@jagr2713 you were on the right track "ok for every 2 servers 4 land out. so if he gets 4 serves 8 land out so he gt 28 serves. So /.... not the relationship was 4/2= and 8/4=2 so what number divided by 2 gives you 28 or what number=2*28

jagr2713 (jagr2713):

oh so i was wrong D;

OpenStudy (freckles):

meant to use word now instead of not

OpenStudy (freckles):

@s117 do you understand this answer ?

jagr2713 (jagr2713):

i apologize @S117

OpenStudy (freckles):

@jagr2713 As I said you were on the right track like you noticed a pattern I don't think you need to apologize This is where the asker could have been like well every lands out number is double the in's number

jagr2713 (jagr2713):

oh ok thanks